# Can Large Language Models Improve Phishing Defense? A Large-Scale Controlled Experiment on Warning Dialogue Explanations

### v1

@@ 4.4 Procedure
The study procedure followed the same one in Greco et al. (2025) and is depicted in Figure 7. To prevent priming effects regarding the experiment’s true objective, participants were initially deceived about the study’s real purpose (Sotirakopoulos et al., 2011; Distler et al., 2021). Participants were told the aim was to evaluate the user experience of a new email client.
@@
After the Prolific task was accepted, the participants were redirected to a web platform where the study was hosted. The (deceptive) purpose of the study, the data usage policies, and the withdrawal right were described on the landing page. According to the ethical guidelines, the participants were asked for their informed digital consent. All the data on participants and interactions were completely anonymized and safely kept on the university servers. Any concerns of the participants could be addressed after the study, and the design was such that it did not cause excessive psychological stress, e.g., phishing scenarios were realistic yet not too alarming.
@@
Once consent was given, all the participants were randomly assigned to four experimental conditions with balanced participation across conditions. Then, participants were requested to follow a case of a fictional user named Alice, who was testing a new email client she had adopted in her workplace. The respondents were asked to interact with the email client by reading messages in the inbox and trying the functionality of embedded links.
@@
The scenario was presented in a banner above the simulated email client interface, which displayed randomised email previews, including the sender’s name, title, and subject (Figure 5). Clicking an email revealed its content; phishing links triggered warnings (Figure 6). After each interaction with a phishing message (via link click or by navigating back), participants answered two attention-check questions: “Who was the sender of the email?” and “Were there links in the email?”. These checks help identify inattentive respondents, which are especially typical in remote settings (Matsuura et al., 2021).
Figure 5: Web platform used for the study.
@@
Figure 6: Warning shown during the study for a fake Amazon email that includes a mimicking URL with a suspicious domain (“.com.cz”).
@@
Upon completing their interactions, participants were debriefed and informed of the actual goal of the study: assessing user responses to warnings triggered by clicks on suspicious links. The debriefing step is a critical ethical practice in studies involving deception (Sotirakopoulos et al., 2011). After the debriefing, participants were asked to reconfirm their consent. Those choosing not to continue would have had their data deleted, though all participants opted to proceed. They then completed a series of questionnaires: first, the warning perception questionnaire from Desolda et al. (2023), followed by the Raw NASA TLX workload assessment (Hart, 1986), a cybersecurity proficiency questionnaire (Olmstead and Smith, 2017), and the Need for Cognition Scale (Coelho et al., 2020). Finally, participants completed a demographic survey stating their age, gender, and average daily internet use. Participants were thanked for their involvement and redirected to Prolific to claim compensation.
@@
Figure 7: Study Procedure - adapted from Greco et al. (2025).
@@ 4.5 Data Analysis
Moreover, to answer RQ3, a deductive thematic analysis (TA) was also conducted to examine qualitative responses from open-ended questions of the warning questionnaire (see items 2, 6, and 9 in Table 1) following Braun and Clarke’s six-step framework (Braun and Clarke, 2006).
@@ 5.1 RQ1 - LLM vs Manual Explanations
To answer the first research question, we compared the results of experimental conditions with the baseline from Greco et al. (2025), which employed warnings with manually generated explanations. To obtain a fair comparison, we excluded from this analysis the conditions that employed counterfactual explanations, as the baseline included a feature-based explanation (n = 300 for the aggregated experimental conditions, n = 150 for the baseline sample, n = 450 in total). The CTRs for the baseline and LLM feature-based conditions (both individual and aggregated) are reported in Figure 8. As it can be observed, LLM-generated warnings generally led to a higher CTR when aggregated, compared to the baseline. A more in-depth analysis revealed that the Claude FB condition had the lowest CTR (9.39%), especially for true positive emails (7.48%), while the baseline led to the lowest CTR in the case of the false positive (9.78%). On the contrary, Llama FB led to the highest CTR in all cases (14.65% in total, 14.62% for true positives, and 14.43% for false positives).
Figure 8: CTRs in the baseline condition (manually-generated, feature-based explanation) and for the LLM-generated feature-based explanations, presented both individually (Claude FB, Llama FB) and aggregated.
The Chi-square test of independence revealed no significant statistical difference between the baseline and the LLM feature-based conditions, both individually and when aggregated ($p>0.05$). This indicates that LLM-generated warnings are at least as effective as manually-generated ones in protecting users. However, comparing the baseline, Claude FB, and Llama FB led to a significant difference ($\chi^{2}=5.86$, df = 2, $p=0.05$) for the True Positive emails. Although pairwise comparisons did not reveal differences, a trend emerged between Claude-generated feature-based warnings (8.90%) and Llama-generated ones (14.29%), as the post-hoc test indicates a marginal difference between the two ($p=0.084$).
@@ 5.2 RQ2 - Feature-based vs Counterfactual Explanations
The second research question regarded the effectiveness of explanations in warning dialogues based on their type (feature-based or counterfactual). Therefore, we aggregated the conditions according to the explanation type they included and made a direct comparison (n=300 for both samples, n=600 in total). To have a balanced analysis, we excluded the baseline from this comparison. The CTRs for the conditions included in this analysis are reported in Figure 9.
Compared to the feature-based explanations, the warnings with counterfactual explanations led to a lower CTR in general (11.41% vs 11.97%) and for the false positive email (10.00% vs 14.06%), while slightly increased the CTR for the true positives (12.04% vs 11.03%).
Figure 9: CTRs for the aggregated conditions with feature-based and counterfactual explanations. The baseline is not included.
@@
To deepen the analysis of RQ2, we investigated the CTR for the experimental conditions taken individually (n=150 for each sample, n=600 in total). Figure 10 reports the results. It is noticeable how feature-based explanations from Claude led to the highest level of protection (CTR = 9.39% in general and CTR = 7.48% for true positives). In contrast, feature-based explanations from Llama led to the highest CTR (14.56% in general, 14.62% for true positives, and 14.43% for the false positives). In the case of the false positive email, the baseline led to the lowest CTR (9.78%).
Figure 10: CTRs for the individual experimental conditions and the baseline.
@@ 5.3 RQ3 - User Perception of Warnings
The third research question explored how users perceive warnings with LLM-generated explanations. To do this, we examined the results of the warning questionnaire (see Table 1) and the raw NASA-TLX. Responses to the NASA-TLX were considered to measure the workload caused by the experimental condition. Moreover, to get more insights, we also computed the six sub-dimensions of the NASA-TLX, i.e., mental demand, physical demand, temporal demand, perceived performance, perceived effort, and frustration level.
@@
Figure 11: Average values of perceived characteristics emerging from the results of the warning questionnaire.
Figure 12: Average values of each dimension of the Raw NASA-TLX and overall scores for each condition and aggregation.
@@
Table 2: Answers to Item 8 of the warning questionnaire by warning condition. Percentages represent the proportion of participants selecting each response option.
Condition “Don’t continue” “Be careful” “Continue” “None”
@@
FB aggregate 76.0% 21.3% 2.0% 0.7%
CF aggregate 72.7% 23.3% 3.0% 1.0%
Claude aggregate 74.7% 22.0% 2.7% 0.7%
Llama aggregate 74.0% 22.7% 2.3% 1.0%
@@ 5.4 RQ4 - How User-Related Factors Affect CTR
To investigate how user-related factors affect CTR, we built a series of binomial logistic regression models including demographic measures (age, gender, expertise, average hours online per day, Need-for-Cognition) and warning-perception variables (understandability, familiarity, trust, perceived suggested action, mental workload). Different models were trained for three outcomes: CTR across all emails (ALL), true positives (TP), and false positives (FP), across all conditions and LLM variants.
Moreover, odds ratios (OR) were computed by exponentiating the estimated coefficients from the logistic regression models, by applying the formula $\mathrm{OR}=e^{\beta}$. This transformation allows interpreting each $\beta$ coefficient as the multiplicative change in the odds of clicking associated with a one-unit increase in the predictor. An OR greater than 1 indicates a higher likelihood of clicking, whereas an OR below 1 indicates a lower probability. To further simplify the interpretation of the OR, we mapped the ORs into effect sizes, categorised as small (OR 1.2–1.5 / 0.67–0.83), moderate (1.51–2 / 0.5–0.66), and large (¿2 / ¡0.5), following the guidelines suggested by Chen et al. Chen et al. (2010). The majority of significant effects observed in our data fell either into the ”small” (OR 1.2–1.3) or ”large” (OR¿2 or ¡0.5) categories, with relatively few predictors yielding odds ratios in the moderate range. Odds ratios of 0.84 to 1.19 were considered as negligible because they show a difference in odds of less than 20% in clicking (which will practically be insignificant even in the case of a significant difference). This strategy is guided by the prescriptions of the applied behavioural research to differentiate between the statistical and practical significance of the requirements (Kirk, 1996; Fritz et al., 2012).
@@
Interactions between being male and need-for-cognition produced small to large effects: OR=1.44 in Counterfactual aggregate for false positives, OR=2.15 in Claude Counterfactual for false positives, OR=1.33 in Claude Counterfactual for all emails, OR=1.32 in Claude Counterfactual for true positives, a smaller effect of OR=1.21 in Counterfactual aggregate for all emails.
@@
Full results, including ORs, p-values, and effect size categorisations, are detailed in Table LABEL:tab:user_factors_ctr reported in A.6.
@@ 5.5 Qualitative results
This subsection describes the themes identified through deductive thematic analysis of Items 2, 6, and 9, as presented in Table 1. This analysis strictly followed Braun and Clarke’s six-step framework (Braun and Clarke, 2006): data familiarisation, code generation, theme identification, theme review, theme definition and naming, and report production. The deductive approach was selected to align the coding framework with the prior study that inspired our experimental design and provided the baseline warning condition (Greco et al., 2025). Two independent researchers with a background in Human-Computer Interaction and qualitative research methods carried out the process. They were blind to the participants’ experimental conditions during the first phases to avoid bias. In particular, the researchers first familiarized themselves with the data by repeatedly reading transcripts and noting initial impressions. Then, they independently coded relevant excerpts. Inter-rater reliability was evaluated using Cohen’s kappa (κ) values: Item 2 = .81, Item 6 = .84, and Item 9 = .79. These values indicate good agreement (Landis and Koch, 1977). Discrepancies were resolved through discussion until a consensus was reached. Related codes were grouped into preliminary themes (80% initial agreement), which were collaboratively refined to ensure an accurate representation of the data. Each theme was then clearly defined and named, and the final report included interpretive commentary and supporting data excerpts. For each of the three items, the themes captured the essence and meaning of the users’ answers, namely, the users’ initial reactions to the warnings (item 2), terms perceived as confusing (item 6), and interpretations of warning meanings (item 9).
@@
Figure 13: Distribution of first reaction themes (Item 2) across experimental conditions, including the manually crafted feature-based baseline from Greco et al. (2025), and the combination of LLM (LLaMA, Claude), explanation style (feature-based, counterfactual). Each theme’s total number of instances is indicated on top of each bar.
@@ 5.5.1 First reaction to warning (Item 2)
In Item 2, participants answered the question: “When you saw the warning, what was your first reaction?”. Five recurring themes were identified across the four experimental conditions: Perceived threat, Take safe action, Further checking, Warning not trusted, and Positive reaction to the warning. These themes are reported in Figure 13, which shows their frequency across the four experimental conditions defined by the LLM (LLaMA or Claude), the explanation style (feature-based or counterfactual), and the manual feature-based explanation from Greco et al. (2025).
@@ 5.5.2 Confusing words (Item 6)
Table 3: Themes, counts, and example answers for confusing words item: Which word/s did you find confusing or too technical?.
Themes Theme count Answer examples
@@
Unclear general word 109 “The numbers provided in the texts, such as the email points” (P52), “The word Deceptive (P41)
@@
Figure 14: Distribution of warning meaning themes (Item 9) across experimental conditions, including the manually crafted feature-based baseline from Greco et al. (2025), and the combination of LLM (LLaMA, Claude), explanation style (feature-based, counterfactual). Each theme’s total number of instances is at the top of each bar.
@@ 5.5.3 Warning meaning (Item 9)
Malicious link. In some instances, participants understood the warning as indicating that the threat was embedded in the link. The perceived level of risk varied, ranging from a clear and immediate danger (e.g., “That the link I clicked leads to a site that doesn’t belong to Facebook and they will steal the information I type.”, P361) to a more cautious interpretation of potential harm (e.g., “Warns users to be careful of suspicious links/emails that can take their data”, P215). This theme was more frequent for counterfactual (26 times in the case of Claude CF and 32 times for Llama CF) and baseline explanations than for feature-based (14 times in the case of Claude FB and 15 times for Llama FB).
Device at risk. The warning was interpreted as an indication that the user’s device could be at risk, for example, due to potential malware contained within the email content or linked website (e.g., “It means the webpage or link you are about to enter may harm your computer of breach data”, P513). This theme was generally less frequent than others, and quite balanced across all conditions.

### v2

@@ 4.4 Procedure
The study procedure followed the same one in greco2025enhancing and is depicted in Figure 7. To prevent priming effects regarding the experiment’s true objective and to mitigate the Hawthorne effect macefield2007usability, participants were initially deceived about the study’s real purpose  (sotirakopoulos2011challenges). Deception is one of the most adopted techniques in cybersecurity usability studies distler2021slr. In our context, explicitly mentioning ”phishing” or ”security” would have biased participants towards a suspicious mindset, unrepresentative of daily email usage. Therefore, participants were told the aim was to evaluate the user experience of a new email client interface. This framing encouraged participants to focus on the functionality and usability of the client, simulating a realistic scenario of split attention where security is a secondary task.
@@
After the Prolific task was accepted, the participants were redirected to a web platform where the study was hosted. The (deceptive) purpose of the study, the data usage policies, and the withdrawal right were described on the landing page. According to the ethical guidelines, participants were asked to provide their informed digital consent. All data on participants and interactions were completely anonymized and securely stored on the university servers. Any concerns that the participants may have had could be addressed after the study. The design was such that it did not cause excessive psychological stress; for example, phishing scenarios were realistic yet not too alarming.
@@
Once consent was obtained, all participants were randomly assigned to four experimental conditions, with balanced participation across conditions. Then, participants were asked to follow a case study of a fictional user named Alice, who was testing a new email client she had adopted at her workplace. The respondents were asked to interact with the email client by reading messages in the inbox and trying the functionality of embedded links.
@@
The scenario was presented in a banner above the simulated email client interface, which displayed randomised email previews, including the sender’s name, title, and subject (Figure 5). Clicking an email revealed its content; phishing links triggered warnings (Figure 6). After each interaction with a phishing message (via link click or by navigating back), participants answered two attention-check questions: “Who was the sender of the email?” and “Were there links in the email?”. These checks help identify inattentive respondents, which are especially typical in remote settings (matsuura2021careless).
Figure 5: Web platform used in the study.
@@
Figure 6: Warning shown during the study for a fake Amazon email containing a mimicking URL with a suspicious domain (“.com.cz”).
@@
Upon completing their interactions, participants were debriefed and informed of the actual goal of the study: assessing user responses to warnings triggered by clicks on suspicious links. The debriefing step is a critical ethical practice in studies involving deception (sotirakopoulos2011challenges). After the debriefing, participants were asked to reconfirm their consent. Those who chose not to continue would have had their data deleted; however, all participants opted to proceed. They then completed a series of questionnaires: first, the warning perception questionnaire from desolda2023explanations, followed by the Raw NASA TLX workload assessment (hart1986NASATLX), a cybersecurity proficiency questionnaire (olmstead2017securityquestionnaire), and the Need for Cognition Scale (coelho2020needforcognition6item). Finally, participants completed a demographic survey that asked about their age, gender, and average daily internet use. Participants were thanked for their involvement and redirected to Prolific to claim compensation.
@@
Figure 7: Study procedure (adapted from greco2025enhancing).
@@ 4.5 Data Analysis
Moreover, to answer RQ3, a deductive thematic analysis (TA) was also conducted to examine qualitative responses from open-ended questions of the warning questionnaire (see items 2, 6, and 9 in Table 1) following Braun and Clarke’s six-step framework (braun2006thematic).
@@ 5.1 RQ1 - LLM vs Manual Explanations
To answer the first research question, we compared the results of the experimental conditions with the baseline from greco2025enhancing, which employed warnings with manually generated explanations. To obtain a fair comparison, we excluded from this analysis conditions that employed counterfactual explanations, as the baseline included a feature-based explanation (n = 300 for the aggregated experimental conditions, n = 150 for the baseline sample, totaling 450). The CTRs for the baseline and LLM feature-based conditions (both individual and aggregated) are reported in Figure 8. As it can be observed, LLM-generated warnings generally led to a higher CTR when aggregated, compared to the baseline. A more in-depth analysis revealed that the Claude FB condition had the lowest CTR (9.39%), especially for true positive emails (7.48%), while the baseline led to the lowest CTR in the case of the false positives (9.78%). On the contrary, Llama FB led to the highest CTR in all cases (14.65% in total, 14.62% for true positives, and 14.43% for false positives).
Figure 8: Click-through rates (CTRs) in the baseline condition (manually generated, feature-based explanation) and for LLM-generated feature-based explanations, shown individually (Claude FB, Llama FB) and aggregated.
The difference in CTR between the Baseline and Claude FB was not statistically significant at the $\alpha=0.05$ level ($p>0.05$). This finding is fundamental: it demonstrates that replacing expensive human experts with an automated LLM (Claude) does not degrade user protection. The LLM’s non-inferiority is a positive result for the scalability of phishing defense. However, comparing the baseline, Claude FB, and Llama FB led to a significant difference ($\chi^{2}=5.86$, df = 2, $p=0.05$) in the number of True Positive emails. Although pairwise comparisons did not reveal differences, a trend emerged between Claude-generated feature-based warnings (8.90%) and Llama-generated ones (14.29%), as the post-hoc test indicates a marginal difference between the two ($p=0.084$).
@@ 5.2 RQ2 - Feature-based vs Counterfactual Explanations
The second research question concerned the effectiveness of explanations in warning dialogues, specifically in terms of their type (feature-based or counterfactual). Therefore, we aggregated the conditions according to the explanation type they included and made a direct comparison (n = 300 for both samples, totaling n = 600). To have a balanced analysis, we excluded the baseline from this comparison. The CTRs for the conditions included in this analysis are reported in Figure 9.
Compared to the feature-based explanations, the warnings with counterfactual explanations led to a lower CTR in general (11.41% vs 11.97%) and for the false positive email (10.00% vs 14.06%), while slightly increasing the CTR for the true positives (12.04% vs 11.03%).
Figure 9: Click-through rates (CTRs) for aggregated conditions with feature-based and counterfactual explanations. The baseline condition is excluded.
@@
To deepen the analysis of RQ2, we investigated the CTR for the experimental conditions taken individually (n = 150 for each sample, totaling n = 600). Figure 10 reports the results. It is noticeable how feature-based explanations from Claude led to the highest level of protection (CTR = 9.39% in general and CTR = 7.48% for true positives). In contrast, feature-based explanations from Llama yielded the highest CTR (14.56% in general, 14.62% for true positives, and 14.43% for false positives). In the case of the false positive email, the baseline resulted in the lowest CTR (9.78%).
Figure 10: Click-through rates (CTRs) for individual experimental conditions and the baseline.
@@ 5.3 RQ3 - User Perception of Warnings
The third research question explored how users perceive warnings with LLM-generated explanations. To achieve this, we examined the results of the warning questionnaire (see Table 1) and the raw NASA-TLX scores. Responses to the NASA-TLX were considered to measure the workload caused by the experimental condition. Moreover, to gain further insights, we also computed the six sub-dimensions of the NASA-TLX, namely mental demand, physical demand, temporal demand, perceived performance, perceived effort, and frustration level.
@@
Figure 11: Average perceived characteristic values derived from the warning questionnaire results.
Figure 12: Average scores for each dimension of the Raw NASA-TLX and overall scores by condition and aggregation.
@@
Table 2: Participants’ actions (Item 8) by warning condition (%).
Condition Don’t continue Be careful Continue None
@@
FB agg. 76.0% 21.3% 2.0% 0.7%
CF agg. 72.7% 23.3% 3.0% 1.0%
Claude agg. 74.7% 22.0% 2.7% 0.7%
Llama agg. 74.0% 22.7% 2.3% 1.0%
@@ 5.4 RQ4 - How User-Related Factors Affect CTR
To investigate how user-related factors affect CTR, we built a series of binomial logistic regression models that included demographic measures (age, gender, expertise, average hours online per day, and Need-for-Cognition) and warning-perception variables (understandability, familiarity, trust, perceived suggested action, and mental workload). Different models were trained for three outcomes: CTR across all emails (ALL), true positives (TP), and false positives (FP), across all conditions and LLM variants.
Moreover, odds ratios (OR) were computed by exponentiating the estimated coefficients from the logistic regression models, by applying the formula $\mathrm{OR}=e^{\beta}$. This transformation allows interpreting each $\beta$ coefficient as the multiplicative change in the odds of clicking associated with a one-unit increase in the predictor. An OR greater than 1 indicates a higher likelihood of clicking, whereas an OR below 1 indicates a lower probability. To further simplify the interpretation of the OR, we mapped the ORs into effect sizes, categorised as small (OR 1.2–1.5 / 0.67–0.83), moderate (1.51–2 / 0.5–0.66), and large (>>2 / <<0.5), following the guidelines suggested in chen2010big. The majority of significant effects observed in our data fell either into the ”small” (OR 1.2–1.3) or ”large” (OR>>2 or <<0.5) categories, with relatively few predictors yielding odds ratios in the moderate range. Odds ratios of 0.84 to 1.19 were considered as negligible because they show a difference in odds of less than 20% in clicking (which will practically be insignificant even in the case of a significant difference). This strategy is guided by the prescriptions of the applied behavioural research to differentiate between the statistical and practical significance of the requirements (Kirk1996; fritz2012effect).
@@
Interactions between being male and need-for-cognition produced small to large effects: OR=1.44 in Counterfactual aggregate for false positives, OR=2.15 in Claude Counterfactual for false positives, OR=1.33 in Claude Counterfactual for all emails, OR=1.32 in Claude Counterfactual for true positives, and a smaller effect of OR=1.21 in Counterfactual aggregate for all emails.
@@
Full results, including ORs, p-values, and effect size categorisations, are detailed in Table LABEL:tab:user_factors_ctr reported in Appendix 11.
@@ 5.5 Qualitative results
This subsection describes the themes identified through deductive thematic analysis of Items 2, 6, and 9, as presented in Table 1. This analysis strictly followed Braun and Clarke’s six-step framework (braun2006thematic): data familiarisation, code generation, theme identification, theme review, theme definition and naming, and report production. The deductive approach was selected to align the coding framework with the prior study that inspired our experimental design and provided the baseline warning condition (greco2025enhancing). Two independent researchers, with backgrounds in Human-Computer Interaction and qualitative research methods, carried out the process. They were blind to the participants’ experimental conditions during the first phases to avoid bias. In particular, the researchers first familiarized themselves with the data by repeatedly reading transcripts and noting initial impressions. Then, they independently coded relevant excerpts. Inter-rater reliability was evaluated using Cohen’s kappa ($\kappa$) values: Item 2 = .81, Item 6 = .84, and Item 9 = .79. These values indicate good agreement (landis1977application). Discrepancies were resolved through discussion until a consensus was reached. Related codes were grouped into preliminary themes (80% initial agreement), which were collaboratively refined to ensure an accurate representation of the data. Each theme was then clearly defined and named, and the final report included interpretive commentary and excerpts of supporting data. For each of the three items, the themes captured the essence and meaning of the users’ answers, specifically their initial reactions to the warnings (item 2), terms perceived as confusing (item 6), and interpretations of warning meanings (item 9).
@@
Figure 13: Distribution of first reaction themes (Item 2) across experimental conditions, including the manually crafted feature-based baseline from greco2025enhancing, and the combination of LLMs (LLaMA, Claude) and explanation styles (feature-based, counterfactual). The total number of instances for each theme is shown atop each bar.
@@ 5.5.1 First reaction to warning (Item 2)
In Item 2, participants answered the question: “When you saw the warning, what was your first reaction?”. Five recurring themes were identified across the four experimental conditions: Perceived threat, Take safe action, Further checking, Warning not trusted, and Positive reaction to the warning. These themes are reported in Figure 13, which shows their frequency across the four experimental conditions defined by the LLM (LLaMA or Claude), the explanation style (feature-based or counterfactual), and the manual feature-based explanation from greco2025enhancing.
@@ 5.5.2 Confusing words (Item 6)
Table 3: Themes, counts, and example answers for confusing words item: Which word(s) did you find confusing or too technical?
Theme Count Example answers
@@
Unclear general word 109 “The numbers provided in the texts, such as the email points” (P52); “The word Deceptive (P41)
@@
Figure 14: Distribution of warning meaning themes (Item 9) across experimental conditions, including the manually crafted feature-based baseline from greco2025enhancing, and the combination of LLMs (LLaMA, Claude) and explanation styles (feature-based, counterfactual). The total number of instances for each theme is displayed at the top of each bar.
@@ 5.5.3 Warning meaning (Item 9)
Malicious link. In some instances, participants understood the warning as indicating that the threat was embedded in the link. The perceived level of risk varied, ranging from a clear and immediate danger (e.g., “That the link I clicked leads to a site that doesn’t belong to Facebook and they will steal the information I type.”, P361) to a more cautious interpretation of potential harm (e.g., “Warns users to be careful of suspicious links/emails that can take their data”, P215). This theme was more frequent in counterfactual (26 times for Claude CF and 32 times for Llama CF) and baseline explanations than in feature-based (14 times for Claude FB and 15 times for Llama FB) explanations.
Device at risk. The warning was interpreted as an indication that the user’s device could be at risk, for example, due to potential malware contained within the email content or linked website (e.g., “It means the webpage or link you are about to enter may harm your computer or breach data”, P513). This theme was generally less frequent than others and quite balanced across all conditions.
@@
5.5.4 Impact of Comprehension on Decision Making
Beyond identifying isolated themes, we cross-referenced these qualitative insights with quantitative behavioral metrics to validate the causal link between explanation content and user action. Notably, users in the Claude condition who avoided the phishing link frequently cited specific details from the generated text (e.g., ”The warning mentioned a mismatch in the URL”) in their open-ended responses. Conversely, users who clicked often dismissed the warning as a ”false alarm” rather than misunderstanding the text. This triangulation confirms that the content of the explanation, and not just the presence of a warning, influenced the decision-making process, validating the intervention’s effectiveness beyond the generic effects of a warning, such as those without explanations.
5.6 Linking Comprehension to Behavior
To address the causality between explanation content and user action, we cross-referenced qualitative feedback with quantitative behavioral metrics. Notably, users in the Claude condition who avoided the phishing link frequently cited specific details from the explanation (e.g., ”The warning mentioned a mismatch in the URL”) in their open-ended responses. Conversely, users who clicked often dismissed the warning as a ”false alarm” rather than misunderstanding the text. This suggests that the content of the explanation directly influenced the decision-making process, validating the intervention’s effectiveness beyond generic warning effects.